### v1

# generalised polynomial chaos

Abstract
This paper presents an approach for the modelling of dependent random variables using generalised polynomial chaos. This allows to write chance-constrained optimization problems with respect to a joint distribution modelling dependencies between different stochastic inputs. Arbitrary dependencies are modelled by using Gaussian copulas to construct the joint distribution. The paper exploits the problem structure and develops suitable transformations to ensure tractability. The proposed method is applied to a probabilistic power reserve procurement problem. The effectiveness of the method to capture dependencies is shown by comparing the approach with a standard approach considering independent random variables.
I Introduction
distributed energy resources are increasingly participating in ancillary service (AS) markets. Many of these resources. such as heat pumps, electric vehicles and photovoltaic (PV) installations have uncertain capabilities. To facilitate the participation of these DERs, some transmission system operators already allow bids for which the full power availability is only guaranteed with a certain probability [1]. This has important consequences for the reliable procurement of reserve capacity. Add short review on approaches for reserve procurement. Look for probabilistic approaches + especially considering dependencies of multiple DER types.
Recently, generalised polynomial chaos (gPC) has been used in power systems applications to model propagation of uncertain input variables. Advantages include the ability to model general probability distributions and represent stochastic variables through orthogonal polynomial expansions. This approach was first used for stochastic power flow computations [4], with inequality constraints formulated only with respect to the expected value. Later work, such as [5], introduces the use of gPC for the solution of optimal power flow problems. Typically, approaches modelling uncertainty with gPC formulate chance constraints through a moment-based reformulation. For Gaussian random variables or distributionally robust (DR) uncertainty, exact reformulations are found [6]. For other distributions, [3] proposes an iterative tuning approach for the bounds in the moment-based reformulation in an attempt to recover the desired constraint satisfaction rate. A major limitation of these approaches, is that they consider the different random input variables to be independent. To allow modelling dependencies between random variables, [7] proposes to construct an orthonormal polynomial basis using a Gram-Schmidt (GS) procedure. A challenge with this approach is that it leads to numerical issues due to the poor conditioning of the GS procedure[2]. Other methods to model dependent random variables with gPC exist, as reported for example in [2]. The Rosenblatt transformation is a mapping that transforms dependent random variables to a space of independent random variables, in which an orthonormal polynomial basis can be constructed using standard approaches. This has for example been applied to solve stochastic power flows [9]. However, this approach requires to determine the inverse Rosenblatt transform for each sample used to obtain polynomial chaos expansion (PCE) coefficients. here further works on dependent gPC?
None of the works above allow to assess the probability of violating the power reserve requirements when considering uncertain bids exhibiting dependencies. Additionally, the proposed approaches to extend gPC to dependent variables, lead to computational issues due to error propagation and computation time. Therefore, this paper makes the following contributions:
  1. 1.
    An approach to model arbitrary dependencies between random variables using (Gaussian) copulas.
  2. 2.
    A tractable computation scheme for the gPC basis, using quadrature integration and exploiting the copula structure.
  3. 3.
    A convex reformulation for probabilistic power reserve procurement considering dependent random variables.
II Methodology
II-A gPC with Dependent Random Variables
Consider the random vector $\bm{\xi}=[\xi_{1},\ldots,\xi_{d}]^{T}$ with joint PDF $p(\bm{\xi})$ and support $\Xi$.
A known method to deal with uncertainty (is the use of a Polynomial Chaos Expansion (PCE), casting a function $f(\bm{\xi})$ of random variables $\bm{\xi}$ into a sum of orthonormal polynomials $\psi(\bm{\xi})$:
$f(\bm{\xi})=\sum_{l=0}^{\infty}a_{l}\psi_{l}(\bm{\xi}),$ (1)
where orthonormality is required with respect to the expectation value
$\mathbb{E}[\psi_{m}(\bm{\xi})\psi_{n}(\bm{\xi})]=\int_{\Xi}\psi_{m}(\bm{\xi})\psi_{n}(\bm{\xi})p(\bm{\xi})d\bm{\xi}=\delta_{mn}.$ (2)
The coefficients can be computed by projection onto the polynomial basis, where the scalar product is identical to the expectation value of the product:
$a_{l}=\left\langle f(\bm{\xi}),\psi_{l}(\bm{\xi})\right\rangle=\mathbb{E}[f(\bm{\xi})\psi_{l}(\bm{\xi})]$ (3)
The main difficulty is the search for orthonormal polynomials with respect to the arbitrary PDF $p(\xi)$. As proven in (XXXSomething), polynomials arising from Askey-scheme are orthogonal for their respective probability distribution (e.g. Hermite polynomials for Gaussian PDFs), however, this assumes independence between the elements of the random vector. In turn, Navarro et al. suggested in [7] that for arbitrarily correlated random variables, orthonormal polynomials can be constructed from a set of linearly independent polynomials using the Gram-Schmidt orthonormalization scheme.
II-B Joint Probability using Copulas
For practical applications, the joint PDF $p(\bm{\xi})$ needs to be approximated in a tractable manner. This requires both a model of the marginal PDFs of each random variable $\xi_{i}$ as well as their dependence. Dependence modelling can be achieved by using a copula $C$, a function mapping the cumulative distribution functions (CDF) $F_{i}(\xi_{i})$ of each marginal to a joint CDF $F(\bm{\xi})$:
$F(\bm{\xi})=C(F_{1}(\xi_{1}),\ldots,F_{d}(\xi_{d}))$ (4)
As the copula takes CDFs as inputs, it effectively decouples the marginal distributions from the dependence structure. According to Sklar’s theorem (XXX), there exists a unique copula for every possible joint CDF. In practice, one can use an empirical copula, as for example in (Pål paper?), which makes no assumptions on the correlation structure or fit a parametric copula. Archimedean copulas such as Gumbel or Clayton copulas, have a single parameter to capture upper and lower tail-dependence respectively. Gaussian, or more generally Student-t copulas, are defined by a correlation matrix $\Sigma$ and degrees of freedom $\nu$, where in the limit $\nu\rightarrow\infty$, the Student-t becomes a Gaussian copula.
For the application reported here, featuring a high dimensionality a Gaussian copula was chosen. This allows for scalability and mathematically convenient expressions (XXX and what?). Setting $F_{i}(\xi_{i})=u_{i}$, it is defined as:
$C_{\bm{\Sigma}}(u_{1},...,u_{d})=\Phi_{\bm{\Sigma}}(\Phi^{-1}(u_{1}),...,\Phi^{-1}(u_{d}))$ (5)
where $\Phi^{-1}(u_{i})$ is the inverse of the normal CDF and $\Phi_{\bm{\Sigma}}(\bm{z})$ the CDF of the multivariate normal with correlation matrix $\bm{\Sigma}$. The joint PDF of the random vector $\bm{\xi}$ is then expressed as
$p(\bm{\xi})=c(\bm{u})*\prod_{i=1}^{d}f_{i}(\xi_{i})$ (6)
where $c(\bm{u})$ is the copula density
$c(\bm{u})=\frac{\partial^{d}C(u_{1},\dots,u_{d})}{\partial u_{1}\,\partial u_{2}\cdots\partial u_{d}}.$ (7)
In the case of a Gaussian copula, this results in
$c_{\Sigma}(\bm{u})=\frac{\phi_{\bm{\Sigma}}\big(\Phi^{-1}(u_{1}),\ldots,\Phi^{-1}(u_{d})\big)}{\prod_{i=1}^{d}\phi\big(\Phi^{-1}(u_{i})\big)},$ (8)
where $\phi_{\bm{\Sigma}}(\bm{x})$ is the PDF of the multivariate normal with correlation matrix $\bm{\Sigma}$} and $\phi(x_{i})$ the normal PDF.
II-C Quadrature Integration
Both for the computation of the orthonormal polynomials using GS and for the projection of expansion coefficients in Eq. (3), integrals of the form
$\int_{\Xi}g(\bm{\xi})\cdot p(\bm{\xi})d\bm{\xi}$ (9)
have to be evaluated numerically, where $g(\bm{\xi})$ is an arbitrary function. Given the high dimensionality of the integral, evaluation with Monte Carlo sampling or importance sampling schemes would be standard, this however leads to relatively high errors. Further, importance sampling schemes are difficult to implement, as the integrands do not follow a constant distribution
To allow accurate evaluation, quadrature integration is used, this however suffers from computational tractability problems if applied to high-dimensional problems. For this reason a scheme reducing the integration dimension is introduced. First, the integration domain needs to be transformed to a space with known quadrature rules. One such space is the unit cube $0,1]^{d}$ over which the Gauss-Legendre rules can be applied. thus let
$u_{i}=F_{i}(\xi_{i})\in[0,1],\qquad i=1,\dots,d.$ (10)
The Jacobian matrix of this transformation is diagonal with entries $\partial u_{i}/\partial\xi_{i}=f_{i}(\xi_{i})$, hence
$d\bm{\xi}=\frac{d\bm{u}}{\prod_{i=1}^{d}f_{i}(\xi_{i})}.$ (11)
Applying the change of variables and substituting in the definition of $p(\bm{\xi})$ from Eq. (6) yields
$\int_{\Xi}g(\bm{\xi})\cdot p(\bm{\xi})d\bm{\xi}=\int_{[0,1]^{d}}f\big(\bm{F}^{-1}(\bm{u})\big)\,c(\bm{u})\,d\bm{u}$ (12)
where $\bm{F}^{-1}(\bm{u})$ is the componentwise inverse CDF. This expression makes no assumptions on the copula used and hence… However, as the copula density $c(\bm{u})$ is in general neither polynomial nor constant, the quadrature will be inexact. In the specific case of Gaussian copulas, a transformation to latent Gaussian space
$z_{i}=\Phi^{-1}(u_{i})\in(-\infty,\infty),\qquad i=1,\dots,d.$ (13)
is helpful. The Jacobian matrix is again diagonal with entries $\partial u_{i}/\partial z_{i}=\phi(z_{i})$, hence
$d\bm{u}=\Big(\prod_{i=1}^{d}\phi(z_{i})\Big)\,d\bm{z}.$ (14)
Applying the change of variables to the copula density and using the definition of the Gaussian copula density from Eq. (8) yields
$\int_{\Xi}g(\bm{\xi})\cdot p(\bm{\xi})d\bm{\xi}=\int_{\mathbb{R}^{d}}g\big(\bm{F}^{-1}(\Phi(\bm{z}))\big)\;\phi_{\Sigma}(\bm{z})\,d\bm{z}.$ (15)
This form allows for Gauss-Hermite quadrature, which expects a polynomial weighted by an uncorrelated normal distribution. Computing the Cholesky factorisation of the correlation matrix $\bm{\Sigma}=\bm{L}\bm{L^{T}}$ allows to “colour“ the correlation onto the quadrature. Letting $\bm{y}_{n}$ and $w_{n}$ be the quadrature nodes and weights. The linear transformation
$\bm{z_{n}}=\bm{L}\bm{y_{n}}$ (16)
induces the correlation. Overall,
$\int_{\Xi}g(\bm{\xi})\,p(\bm{\xi})\,d\bm{x}\;\approx\;\sum_{n=1}^{N}w_{n}\cdot g(\bm{F}^{-1}(\Phi(\bm{Ly_{n}})))$ (17)
Standard Gauss-Hermite quadrature uses a tensor-product grid, meaning evaluating $k$ points per dimension for a $d_{I}$ dimensional integral leads to $N=k^{d_{I}}$ nodes. The exponential scaling with the integral dimension requires a strict limitation.
II-C1 Numerical Improvements
Two ideas allow for decent computational performance when computing the gPC basis. First, GS sequentially evaluates inner products of polynomials to construct the basis polynomials. As noted in [8], the inner product of two arbitrary polynomials $P_{m}(\bm{\xi}),P_{n}(\bm{\xi})$ constructed from the multivariate monomial set $\{1,\xi_{1},\ldots,\xi_{d}^{\nu}\}$, can be expanded to a sum of monomial expectation values
$\left\langle P_{m}(\bm{\xi}),P_{n}(\bm{\xi})\right\rangle$ $=p_{0}\left\langle 1,1\right\rangle+p_{1}\left\langle 1,\xi_{1}\right\rangle+\ldots+p_{P}\left\langle\xi_{d}^{\nu},\xi_{d}^{\nu}\right\rangle$
$=e_{0}+e_{1}\mathbb{E}[\xi_{1}]+\ldots+e_{E}\mathbb{E}[\xi_{d}^{2\nu}]$ (18)
where $p_{i}$ and $e_{i}$ are the appropriate coefficients. Evaluating the monomial expectation values before applying GS can be done in a parallel fashion, allowing for increased speed.
Second, the expectation values of monomials only depends on dimensions appearing in the monomial, for instance
$\mathbb{E}[\xi_{1}\xi_{2}]=\int_{\Xi_{1,2}}\xi_{1}\xi_{2}\,p_{1,2}(\xi_{1},\xi_{2})\,d\bm{\xi}_{1,2}.$ (19)
For elliptic copulas the reduced PDFs have identical structure to the full joint PDF, so Eq. (6) still holds, but with less input variables.
These two observations limit the maximum integral dimension to be evaluated to $2\nu$ with expectations of the form $\mathbb{E}[\xi_{1}\xi_{2}\ldots\xi_{2\nu}]$. Crucially, this decouples integral dimension $d_{I}$ from random variable dimension $d$, avoiding the curse of dimensionality.
III Use Case
IV Conclusion

### v2

@@
 This has important consequences for the reliable procurement of reserve capacity.
-Add short review on approaches for reserve procurement.
-Look for probabilistic approaches + especially considering dependencies of multiple DER types.
+This raises new questions with respect to the amount of reserve that should be procured, as the availability of the procured reserve becomes a stochastic quantity.
 Recently, generalised polynomial chaos (gPC) has been used in power systems applications to model propagation of uncertain input variables.
@@
 However, this approach requires to determine the inverse Rosenblatt transform for each sample used to obtain polynomial chaos expansion (PCE) coefficients.
-here further works on dependent gPC?
-None of the works above allow to assess the probability of violating the power reserve requirements when considering uncertain bids exhibiting dependencies.
+For linear dependencies, this can be simplified to the Nataf transform [4], but the authors of [2] show that it performs worse than their proposed GS procedure.
+Other approaches, so-called domination methods, approximate functions of dependent random variables with functions of independent random variables, but only show good performance when an accurate approximation can be obtained [2].
+None of the works above handling reserve procurement allow to assess the probability of violating the power reserve requirements when considering uncertain bids exhibiting dependencies.
 Additionally, the proposed approaches to extend gPC to dependent variables, lead to computational issues due to error propagation and computation time.
@@
 II-A gPC with Dependent Random Variables
-Consider the random vector $\bm{\xi}=[\xi_{1},\ldots,\xi_{d}]^{T}$ with joint PDF $p(\bm{\xi})$ and support $\Xi$.
-A known method to deal with uncertainty (is the use of a Polynomial Chaos Expansion (PCE), casting a function $f(\bm{\xi})$ of random variables $\bm{\xi}$ into a sum of orthonormal polynomials $\psi(\bm{\xi})$:
-$f(\bm{\xi})=\sum_{l=0}^{\infty}a_{l}\psi_{l}(\bm{\xi}),$ (1)
-where orthonormality is required with respect to the expectation value
-$\mathbb{E}[\psi_{m}(\bm{\xi})\psi_{n}(\bm{\xi})]=\int_{\Xi}\psi_{m}(\bm{\xi})\psi_{n}(\bm{\xi})p(\bm{\xi})d\bm{\xi}=\delta_{mn}.$ (2)
-The coefficients can be computed by projection onto the polynomial basis, where the scalar product is identical to the expectation value of the product:
-$a_{l}=\left\langle f(\bm{\xi}),\psi_{l}(\bm{\xi})\right\rangle=\mathbb{E}[f(\bm{\xi})\psi_{l}(\bm{\xi})]$ (3)
-The main difficulty is the search for orthonormal polynomials with respect to the arbitrary PDF $p(\xi)$.
-As proven in (XXXSomething), polynomials arising from Askey-scheme are orthogonal for their respective probability distribution (e.g.
-Hermite polynomials for Gaussian PDFs), however, this assumes independence between the elements of the random vector.
-In turn, Navarro et al.
-suggested in [7] that for arbitrarily correlated random variables, orthonormal polynomials can be constructed from a set of linearly independent polynomials using the Gram-Schmidt orthonormalization scheme.
+Consider the random vector $\bm{\xi}=[\xi_{1},\ldots,\xi_{d}]^{\top}$ of dimension $d$ with joint probability density function (PDF) $p(\bm{\xi})$ and support $\Xi$.
+As shown in previous works (e.g.
+[8]), a square-integrable random function $q(\bm{\xi})$ can be represented by a PCE
+$q(\bm{\xi})\cong\sum_{l=0}^{L}a_{l}\psi_{l}(\bm{\xi}).$ (1)
+where the equality becomes exact as the arbitrary number of terms $L$ goes to infinity. The basis polynomials $\{\psi_{l}\}$ are orthonormal with respect to the inner product
+$\left\langle\psi_{m},\psi_{n}\right\rangle=\int_{\Xi}\psi_{m}(\bm{\xi})\,\psi_{n}(\bm{\xi})\,p(\bm{\xi})\,\mathrm{d}\bm{\xi}.$ (2)
+The expansion coefficients $\{a_{l}\}$ are obtained by projection onto the polynomial basis. Since this inner product corresponds to an expectation, the coefficients read
+$a_{l}=\left\langle q,\psi_{l}\right\rangle=\mathbb{E}\!\left[q(\bm{\xi})\,\psi_{l}(\bm{\xi})\right].$ (3)
+The main challenge lies in the construction of orthonormal polynomial bases with respect to an arbitrary joint PDF $p(\bm{\xi})$.
+For independent random variables, orthogonal polynomial families associated with the Askey scheme can be employed, since they are orthogonal with respect to their corresponding marginal distributions (e.g.
+Hermite polynomials for Gaussian random variables).
+For general dependent random variables, the basis can be constructed from a suitably chosen set of multivariate monomials of the random variables (e.g.
+all monomials with degree $\leq\nu$).
+As noted, a GS procedure can be used for the construction, but is numerically ill-conditioned [2].
+To address this, [10] provides an alternative algorithm based on a Whitening transformation, relying only in part on GS procedures.
+Finally, the moments of the output distribution of $q(\bm{\xi})$ can be recovered from the expansion coefficients. For instance the mean and variance are given by
+$\mu$ $=a_{0}$ (4)
+$\sigma$ $\cong\left\|[a_{1},\ldots,a_{L}]^{\top}\right\|_{2}$
 II-B Joint Probability using Copulas
-For practical applications, the joint PDF $p(\bm{\xi})$ needs to be approximated in a tractable manner.
-This requires both a model of the marginal PDFs of each random variable $\xi_{i}$ as well as their dependence.
-Dependence modelling can be achieved by using a copula $C$, a function mapping the cumulative distribution functions (CDF) $F_{i}(\xi_{i})$ of each marginal to a joint CDF $F(\bm{\xi})$:
-$F(\bm{\xi})=C(F_{1}(\xi_{1}),\ldots,F_{d}(\xi_{d}))$ (4)
-As the copula takes CDFs as inputs, it effectively decouples the marginal distributions from the dependence structure.
-According to Sklar’s theorem (XXX), there exists a unique copula for every possible joint CDF.
-In practice, one can use an empirical copula, as for example in (Pål paper?), which makes no assumptions on the correlation structure or fit a parametric copula.
-Archimedean copulas such as Gumbel or Clayton copulas, have a single parameter to capture upper and lower tail-dependence respectively.
-Gaussian, or more generally Student-t copulas, are defined by a correlation matrix $\Sigma$ and degrees of freedom $\nu$, where in the limit $\nu\rightarrow\infty$, the Student-t becomes a Gaussian copula.
-For the application reported here, featuring a high dimensionality a Gaussian copula was chosen.
-This allows for scalability and mathematically convenient expressions (XXX and what?).
-Setting $F_{i}(\xi_{i})=u_{i}$, it is defined as:
-$C_{\bm{\Sigma}}(u_{1},...,u_{d})=\Phi_{\bm{\Sigma}}(\Phi^{-1}(u_{1}),...,\Phi^{-1}(u_{d}))$ (5)
-where $\Phi^{-1}(u_{i})$ is the inverse of the normal CDF and $\Phi_{\bm{\Sigma}}(\bm{z})$ the CDF of the multivariate normal with correlation matrix $\bm{\Sigma}$.
-The joint PDF of the random vector $\bm{\xi}$ is then expressed as
+For practical modelling of statistically dependent random variables, the joint PDF $p(\bm{\xi})$ must be approximated in a tractable manner.
+This requires specifying both the marginal distributions of the individual components $f(\xi_{i})$ as well as their dependence structure.
+Copulas provide a general and flexible framework for separating the modelling of marginal distributions from the modelling of statistical dependence.
+Specifically, a Copula $C$ links the marginal cumulative distribution functions (CDFs) $F_{i}(\xi_{i})$ to the joint CDF $F(\bm{\xi})$.
+This link is implicitly expressed as
+$F(\bm{\xi})=C\!\left(F_{1}(\xi_{1}),\ldots,F_{d}(\xi_{d})\right).$ (5)
+Sklar’s Theorem states, that for every multivariate CDF there exists a unique Copula capturing the general dependence structure given by (5).
+In practice, one of several parametric Copula families is used.
+For the application considered in this paper no specific tail-dependence is expected.
+As a result, a Gaussian Copula is selected to maintain computational tractability and leverage its closed-form mathematical expressions.
+Identifying $u_{i}=F_{i}(\xi_{i})$, it is defined as:
+$C_{\bm{\Sigma}}(\bm{u})=\Phi_{\bm{\Sigma}}(\Phi^{-1}(u_{1}),...,\Phi^{-1}(u_{d}))$ (6)
+where $\Phi_{\bm{\Sigma}}(\bm{z})$ is the multivariate normal CDF with correlation matrix $\bm{\Sigma}$ and $\Phi^{-1}(u_{i})$ the inverse normal CDF. The joint PDF is then expressed as
 $p(\bm{\xi})=c(\bm{u})*\prod_{i=1}^{d}f_{i}(\xi_{i})$ (7)
-where $c(\bm{u})$ is the copula density
-$c(\bm{u})=\frac{\partial^{d}C(u_{1},\dots,u_{d})}{\partial u_{1}\,\partial u_{2}\cdots\partial u_{d}}.$ (7)
-In the case of a Gaussian copula, this results in
-$c_{\Sigma}(\bm{u})=\frac{\phi_{\bm{\Sigma}}\big(\Phi^{-1}(u_{1}),\ldots,\Phi^{-1}(u_{d})\big)}{\prod_{i=1}^{d}\phi\big(\Phi^{-1}(u_{i})\big)},$ (8)
-where $\phi_{\bm{\Sigma}}(\bm{x})$ is the PDF of the multivariate normal with correlation matrix $\bm{\Sigma}$} and $\phi(x_{i})$ the normal PDF.
+where $c(\bm{u})$ is the Copula density. For a Gaussian Copula, the density is given by
+$c_{\bm{\Sigma}}(\bm{u})=\frac{\phi_{\bm{\Sigma}}\big(\Phi^{-1}(u_{1}),\ldots,\Phi^{-1}(u_{d})\big)}{\prod_{i=1}^{d}\phi\big(\Phi^{-1}(u_{i})\big)},$ (8)
+where $\phi_{\bm{\Sigma}}(\bm{x})$ is the multivariate normal PDF with correlation matrix $\bm{\Sigma}$} and $\phi(x_{i})$ the normal PDF.
 II-C Quadrature Integration
-Both for the computation of the orthonormal polynomials using GS and for the projection of expansion coefficients in Eq. (3), integrals of the form
+Both the construction of the orthonormal polynomial basis and the computation of the expansion coefficients require the evaluation of a large number of integrals of the form
 $\int_{\Xi}g(\bm{\xi})\cdot p(\bm{\xi})d\bm{\xi}$ (9)
-have to be evaluated numerically, where $g(\bm{\xi})$ is an arbitrary function.
-Given the high dimensionality of the integral, evaluation with Monte Carlo sampling or importance sampling schemes would be standard, this however leads to relatively high errors.
-Further, importance sampling schemes are difficult to implement, as the integrands do not follow a constant distribution
-To allow accurate evaluation, quadrature integration is used, this however suffers from computational tractability problems if applied to high-dimensional problems.
-For this reason a scheme reducing the integration dimension is introduced.
-First, the integration domain needs to be transformed to a space with known quadrature rules.
-One such space is the unit cube $0,1]^{d}$ over which the Gauss-Legendre rules can be applied.
-thus let
+with $g(\bm{\xi})$ an arbitrary function.
+In general, these integrals cannot be evaluated analytically and must be approximated numerically.
+Given the potentially high dimensionality of the integrals, Monte Carlo sampling is a natural choice.
+This however requires a large amount of samples to achieve sufficient accuracy and sampling errors may accumulate throughout the PCE process.
+Moreover, the use of importance sampling is not straightforward in this context, since the integrands encountered vary across polynomial orders and random functions.
+Therefore, to enable fast and accurate numerical integration, quadrature integration schemes in combination with dimensionality-reduction strategies are used.
+First, the integration domain is mapped to a space with known quadrature rules.
+One such space is the unit cube $0,1]^{d}$ on which tensor-grid Gauss-Legendre quadrature can be applied.
+Let
 $u_{i}=F_{i}(\xi_{i})\in[0,1],\qquad i=1,\dots,d.$ (10)
-The Jacobian matrix of this transformation is diagonal with entries $\partial u_{i}/\partial\xi_{i}=f_{i}(\xi_{i})$, hence
+The Jacobian matrix of this transformation is diagonal with entries $\partial u_{i}/\partial\xi_{i}=f_{i}(\xi_{i})$ the marginal PDFs. Hence
 $d\bm{\xi}=\frac{d\bm{u}}{\prod_{i=1}^{d}f_{i}(\xi_{i})}.$ (11)
-Applying the change of variables and substituting in the definition of $p(\bm{\xi})$ from Eq. (6) yields
-$\int_{\Xi}g(\bm{\xi})\cdot p(\bm{\xi})d\bm{\xi}=\int_{[0,1]^{d}}f\big(\bm{F}^{-1}(\bm{u})\big)\,c(\bm{u})\,d\bm{u}$ (12)
-where $\bm{F}^{-1}(\bm{u})$ is the componentwise inverse CDF.
-This expression makes no assumptions on the copula used and hence… However, as the copula density $c(\bm{u})$ is in general neither polynomial nor constant, the quadrature will be inexact.
-In the specific case of Gaussian copulas, a transformation to latent Gaussian space
-$z_{i}=\Phi^{-1}(u_{i})\in(-\infty,\infty),\qquad i=1,\dots,d.$ (13)
-is helpful. The Jacobian matrix is again diagonal with entries $\partial u_{i}/\partial z_{i}=\phi(z_{i})$, hence
-$d\bm{u}=\Big(\prod_{i=1}^{d}\phi(z_{i})\Big)\,d\bm{z}.$ (14)
-Applying the change of variables to the copula density and using the definition of the Gaussian copula density from Eq. (8) yields
-$\int_{\Xi}g(\bm{\xi})\cdot p(\bm{\xi})d\bm{\xi}=\int_{\mathbb{R}^{d}}g\big(\bm{F}^{-1}(\Phi(\bm{z}))\big)\;\phi_{\Sigma}(\bm{z})\,d\bm{z}.$ (15)
-This form allows for Gauss-Hermite quadrature, which expects a polynomial weighted by an uncorrelated normal distribution.
-Computing the Cholesky factorisation of the correlation matrix $\bm{\Sigma}=\bm{L}\bm{L^{T}}$ allows to “colour“ the correlation onto the quadrature.
-Letting $\bm{y}_{n}$ and $w_{n}$ be the quadrature nodes and weights.
-The linear transformation
-$\bm{z_{n}}=\bm{L}\bm{y_{n}}$ (16)
-induces the correlation. Overall,
-$\int_{\Xi}g(\bm{\xi})\,p(\bm{\xi})\,d\bm{x}\;\approx\;\sum_{n=1}^{N}w_{n}\cdot g(\bm{F}^{-1}(\Phi(\bm{Ly_{n}})))$ (17)
-Standard Gauss-Hermite quadrature uses a tensor-product grid, meaning evaluating $k$ points per dimension for a $d_{I}$ dimensional integral leads to $N=k^{d_{I}}$ nodes.
-The exponential scaling with the integral dimension requires a strict limitation.
+Applying the change of variables and substituting $p(\bm{\xi})$ from (7) yields
+$\int_{\Xi}g(\bm{\xi})\cdot p(\bm{\xi})d\bm{\xi}=\int_{[0,1]^{d}}g\big(\bm{F}^{-1}(\bm{u})\big)\,c(\bm{u})\,d\bm{u}$ (12)
+where $\bm{F}^{-1}(\bm{u})$ is short-hand notation for the componentwise inverse marginal CDF.
+Evidently, this transformation is valid no matter the Copula used, in principle allowing any Copula family to be used for this quadrature integration scheme.
+However, the Copula density $c(\bm{u})$ is in general neither polynomial nor constant, meaning Gauss-Legendre quadrature will be inexact.
+For the specific case of Gaussian Copulas, consider a second transformation to a Gaussian latent space
+$z_{i}=\Phi^{-1}(u_{i})\in(-\infty,\infty),\qquad i=1,\dots,d$ (13)
+on which tensor-grid Gauss-Hermite quadrature is applicable. The Jacobian matrix is again diagonal with entries $\partial u_{i}/\partial z_{i}=\phi(z_{i})$, hence
+$d\bm{u}=\Big(\prod_{i=1}^{d}\phi(z_{i})\Big)\,d\bm{z}$ (14)
+Applying the change of variables and substituting the Gaussian Copula density from (8) yields
+$\int_{\Xi}g(\bm{\xi})\cdot p(\bm{\xi})d\bm{\xi}=\int_{\mathbb{R}^{d}}g\big(\bm{F}^{-1}(\bm{\Phi}(\bm{z}))\big)\;\phi_{\bm{\Sigma}}(\bm{z})\,d\bm{z}$ (15)
+with $\bm{\Phi}(\bm{z})$ the componentwise normal CDF.
+The result in (15) features a correlated normal PDF.
+However, Gauss-Hermite quadrature expects to weighted by an uncorrelated normal PDF.
+To address this, let $\bm{\Sigma}=\bm{L}\bm{L^{T}}$ be the Cholesky factorisation of the correlation matrix.
+For a given set of quadrature nodes $\{\bm{z}_{n}\}$ and weights $\{w_{n}\}$ the linear transformation
+$\bm{y}_{n}=\bm{L}\bm{z}_{n}$ (16)
+allows to “colour“ the correlation onto the quadrature nodes resulting in
+$\int_{\Xi}g(\bm{\xi})\,p(\bm{\xi})\,d\bm{x}\;\approx\;\sum_{n=1}^{N}w_{n}\cdot g(\bm{F}^{-1}(\Phi(\bm{Lz}_{n})))$ (17)
+The number of nodes is given by $N=k^{d_{I}}$ with $k$ the number of evaluation points per dimension and $d_{I}$ the integral dimensionality.
+Due to this exponential scaling, dimensionality reduction strategies are necessary.
+Note that while the second transformation eliminated the non-polynomial Copula density $c(\bm{u})$, improving numerical accuracy, there is still no exactness guarantee.
+Gauss-Hermite quadrature is exact for polynomials up to degree 2k12k-1 weighted by a normal PDF, but $g(\bm{F}^{-1}(\bm{\Phi}(\bm{Lz})))$ does not necessarily behave like a polynomial of degree $\leq 2k-1$ in Gaussian latent space.
 II-C1 Numerical Improvements
-Two ideas allow for decent computational performance when computing the gPC basis.
-First, GS sequentially evaluates inner products of polynomials to construct the basis polynomials.
-As noted in [8], the inner product of two arbitrary polynomials $P_{m}(\bm{\xi}),P_{n}(\bm{\xi})$ constructed from the multivariate monomial set $\{1,\xi_{1},\ldots,\xi_{d}^{\nu}\}$, can be expanded to a sum of monomial expectation values
-$\left\langle P_{m}(\bm{\xi}),P_{n}(\bm{\xi})\right\rangle$ $=p_{0}\left\langle 1,1\right\rangle+p_{1}\left\langle 1,\xi_{1}\right\rangle+\ldots+p_{P}\left\langle\xi_{d}^{\nu},\xi_{d}^{\nu}\right\rangle$
+The dimensionality-reduction strategy used to make quadrature integration tractable stems from two key observations: First, as noted in [10], inner products between polynomials can be expanded to a sum of monomial expectation values.
+For instance, let $P_{m}(\bm{\xi}),P_{n}(\bm{\xi})$ be two arbitrary polynomials constructed from the set of multivariate monomials of degree $\leq\nu$ $\{1,\xi_{1},\ldots,\xi_{d}^{\nu}\}$.
+Then the inner product
+$\left\langle P_{m},P_{n}\right\rangle$ $=d_{0}\!\left\langle 1,1\right\rangle+d_{1}\!\left\langle 1,\xi_{1}\right\rangle+\ldots+d_{D}\!\left\langle\xi_{d}^{\nu},\xi_{d}^{\nu}\right\rangle$
 $=e_{0}+e_{1}\mathbb{E}[\xi_{1}]+\ldots+e_{E}\mathbb{E}[\xi_{d}^{2\nu}]$ (18)
-where $p_{i}$ and $e_{i}$ are the appropriate coefficients. Evaluating the monomial expectation values before applying GS can be done in a parallel fashion, allowing for increased speed.
-Second, the expectation values of monomials only depends on dimensions appearing in the monomial, for instance
+where $d_{i},e_{i}$ are the appropriate coefficients. Second, the expectation values of monomials only depend on dimensions appearing in the monomial, for instance
 $\mathbb{E}[\xi_{1}\xi_{2}]=\int_{\Xi_{1,2}}\xi_{1}\xi_{2}\,p_{1,2}(\xi_{1},\xi_{2})\,d\bm{\xi}_{1,2}.$ (19)
-For elliptic copulas the reduced PDFs have identical structure to the full joint PDF, so Eq. (6) still holds, but with less input variables.
-These two observations limit the maximum integral dimension to be evaluated to $2\nu$ with expectations of the form $\mathbb{E}[\xi_{1}\xi_{2}\ldots\xi_{2\nu}]$.
-Crucially, this decouples integral dimension $d_{I}$ from random variable dimension $d$, avoiding the curse of dimensionality.
+where $p_{1,2}(\xi_{1},\xi_{2})$ is a reduced joint PDF.
+For elliptic Copulas they have identical structure to the full joint PDF defined in (7), just with less input variables.
+For other Copulas, the reduced joint PDFS might need to be retrieved separately.
+Now, instead of directly applying an orthonormalization scheme, the monomial expectation values of all pairs of monomials in the chosen set can be evaluated beforehand.
+When applying the orthonormalization scheme, the inner products can be expanded as in (18) and the expectation values substituted.
+Crucially, precomputing expectation values can be done in a parallel fashion and naturally limits the maximum integral dimension to be evaluated to $2\nu$.
+This because the highest dimensional expectation values to be evaluated are of the form $\mathbb{E}[\xi_{1}\xi_{2}\ldots\xi_{2\nu}]$.
+This approach decouples the integral dimension $d_{I}$ from the random variable dimension $d$.
+In the case where multiple random functions $q_{i}(\bm{\xi_{i}})$ only depending on subsets $\bm{\xi_{i}}$ of the random vector are expanded to the same polynomial basis, a similar approach can be used for the expansion coefficients.
+Here the expectation values $\mathbb{E}[q_{i}(\bm{\xi_{i}})],\ldots,\mathbb{E}[q_{i}(\bm{\xi_{i}})\xi_{d}^{\nu}]$ can be computed before with reduced dimension.
+Note, that the reduction in this case is dependent on the dimensionality of the random functions $q_{i}(\bm{\xi_{i}})$.
+Finally, the set of monomials from which the polynomial basis is constructed can be tuned.
+For instance, assume the random functions $q_{i}(\bm{\xi_{i}})$ are models of PV power output at locations $i$ with $\bm{\xi_{i}}$ the local weather conditions like irradiation and temperature.
+Physically meaningless monomials such as the product of squares of the irradiations at two different locations can be dropped from the set without significantly affecting the expansion accuracy.
 III Use Case
+Reserve power, procured from AS markets, is activated whenever the grid frequency exceeds or dips below certain thresholds to ensure stable frequency levels.
+Given the critical nature of reserve activation, the risk of under-delivery must be kept to a minimum.
+This need for security can lead to high procurement costs.
+A recent initiative to keep economic efficiency is the Allocation of Cross-zonal Capacity and Procurement of aFRR Cooperation Agreement (ALPACA) which co-optimizes the procurement of reserves across borders of Austria, Czechia and Germany.
+To validate the formalism introduced in this paper, a simple cross-border joint procurement optimization with uncertain bids and chance constraints on the available power is solved.
+Formally, let $X$ and $Y$ be two zones sharing a common market with a set of bids $E_{i}(\bm{\xi})$ and costs $c_{i}$.
+Bids $i={1,\ldots,N_{X}}$ are located in zone $X$ and $i={N_{X}+1,\ldots,N_{X}+N_{Y}}$ are located zone Y.
+The decision variables $x_{i},y_{i}\in[0,1]$ determine the procured share of each bid.
+The objective is to minimize total cost, while both zones meet their reserve needs $R^{(X)},R^{(Y)}$ and stay below exchange limits $T_{X\to Y},T_{Y\to X}$ with $99\%$ probability.
+Expressed in standard form
+$\min_{\bm{x},\bm{y}}$ $\bm{c^{T}}(\bm{x}+\bm{y})$ (20)
+s.t. $0\leq\bm{x}+\bm{y}\leq 1$
+[…]
+[…]
+[…]
+[…]
+with the first constraint ensuring each bid is procured at most once.
+To solve this problem, the chance-constraints have to be reformulated.
+For this, note that arbitrary weighted sums of gPC expansions result in a new gPC expansion
+$\sum_{j=1}^{Z}z_{j}E_{j}(\bm{\xi})$ $=\sum_{j=1}^{Z}z_{j}\sum_{l=0}^{L}a_{l}^{j}\psi_{l}(\bm{\xi})$ (21)
+$=\sum_{l=0}^{L}\bar{a}_{l}\psi_{l}(\bm{\xi})$
+Figure 1: Illustration of the use case with four bids per zone. Two bids each are distributed across four locations with each two local random variables describing the temperature and irradiation.
+where $\bar{a}_{l}=\sum_{j=1}^{Z}z_{j}a_{l}^{j}$. Using the properties in (4), the mean and variance of such a weighted sum can be expressed as functions of the weights
+$\mu$ $=\sum_{j=1}^{Z}z_{j}a_{0}^{j}$ (22)
+$\sigma$ $=\left\|\begin{bmatrix}a_{1}^{1}&\ldots&a_{1}^{Z}\\ \vdots&\ddots&\vdots\\ a_{L}^{1}&\ldots&a_{L}^{Z}\end{bmatrix}\begin{bmatrix}z_{1}\\ \vdots\\ z_{Z}\end{bmatrix}\right\|_{2}=\left\|(\bm{Az})_{Z}^{1}\right\|_{2}$
+where $(\cdot)^{1}_{Z}$ indicates the index range of the matrix columns and vector. This allows a convex reformulation of the chance-constraints using quantile approximation, for instance
+[…]
+$\sum_{i=1}^{N_{X}+N_{Y}}x_{i}a_{0}^{i}+\lambda_{1}\left\|(\bm{Ax})^{1}_{N_{X}+N_{Y}}\right\|_{2}\geq R^{(X)}$
+where $\lambda_{1}$ is the quantile factor for which $\mu+\lambda_{1}*\sigma$ corresponds to the $1^{st}$ percentile of the distribution.
+For normal distributions $\lambda_{1},\lambda_{99}=\mp 2.326$.
+Applying this reformulation to all chance-constraints in problem (20) results in a convex optimization problem.
+$\min_{\bm{x},\bm{y}}$ $\bm{c^{T}}(\bm{x}+\bm{y})$ (24)
+s.t. $0\leq\bm{x}+\bm{y}\leq 1$
+$\sum_{i=1}^{N_{X}+N_{Y}}x_{i}a_{0}^{i}+\lambda_{1}\left\|(\bm{Ax})_{N_{X}+N_{Y}}^{1}\right\|_{2}\geq R^{(X)}$
+$\sum_{i=1}^{N_{X}+N_{Y}}y_{i}a_{0}^{i}+\lambda_{1}\left\|(\bm{Ay})_{N_{X}+N_{Y}}^{1}\right\|_{2}\geq R^{(Y)}$
+$\;\;\;\sum_{i=1}^{N_{X}}\;\;\;y_{i}a_{0}^{i}+\lambda_{99}\left\|(\bm{Ay})_{N_{X}}^{1}\right\|_{2}\leq T_{X\to Y}$
+$\sum_{i=N_{X}+1}^{N_{X}+N_{Y}}x_{i}a_{0}^{i}+\lambda_{99}\left\|(\bm{Ax})_{N_{X}+N_{Y}}^{N_{X}+1}\right\|_{2}\leq T_{Y\to X}$
+We solved a concrete example with $N_{X},N_{Y}=4$ bids per zone and $d=8$ random variables.
+In two situations.
+The bid functions are assumed to represent PV and HP capabilities.
+The random variables stand for the temperature and irradiation at four different locations with each two bids.
+Fig.
+1 illustrates the setup.
+III-1 Normal marginals
+As $\lambda$ values are known exactly for normal distributions, the full process is first validated with arbitrary normally distributed marginals and linear bid functions.
+The correlation matrix $\bm{\Sigma}$ is fit from synthetic weather data of four different locations retrieved from renewables.ninja [9], however as there is no validation with real data, an arbitrary $\bm{\Sigma}$ could be used.
+The polynomial basis is constructed from 9 monomials (all monomials of degree $\leq 1$).
+The quality of the gPC expansion is checked by drawing random samples from the Copula and evaluating both the bid functions and the expansion.
+The resulting distributions are displayed in Fig.
+2.
+As constructed, the bid power distributions are indeed normally distributed.
+Moreover, the gPC expansion accurately approximates the bid functions.
+The maximum numerical error of the expansion was found to be on the order of $10^{-7}\%$
+Figure 2: Sampled bid function distributions computed both with the bid functions and gPC expansions for the eight bids in the case of normally distributed marginals and linear bid functions.
+Problem (24) is then solved with parameters $R^{(X)}$, $R^{(Y)}$, $T_{X\to Y}$, $T_{Y\to X}=100$ and $c_{1-8}=1$ to obtain optimal $\bm{x},\bm{y}$.
+The chance-constraints in (20) are validated by sampling both from the Copula and from the marginals in an uncorrelated.
+The resulting distributions are shown in Fig.
+3.
+As expected, the optimization problem correctly satisfies the chance-constraints only for the correlated case.
+This results suggests, that the presented formalism works as intended.
+Figure 3: Sampled procured power distributions computed with the gPC expansion in the case of normally distributed marginals.
+For comparison the PDFs both considering and not considering correlation were sampled.
+The empirical $1^{st}$ and $99^{th}$ percentiles are indicated for verification of the chance-constraints.
+III-2 Arbitrary marginals
+Obviously, bid functions are in general not linear in the random variables and the marginals not normally distributed.
+For example, solar irradiance uncertainty is frequently modelled using Beta distributions due to its bounded support.
+For validation in a more realistic setting, the bid functions are replaced by simple polynomial models of PV and HP capability and the marginals by representative Beta distributions.
+The polynomial basis was constructed from 21 physically relevant monomials.
+The computation of the basis and all coefficients was achieved in roughly 2 minutes total with a non-parallelised implementation and $k=15$ quadrature points per dimension.
+Validating the gPC expansions with sampling resulted in the bid power distributions shown in Fig.
+4.
+The maximum expansion error was found to be on the order of $10^{-6}\%$
+Problem (24) was solved with parameters $R^{(X)}$, $R^{(Y)}=1000$, $T_{X\to Y}$, $T_{Y\to X}=500$ and $c_{1-8}=1$.
+The constraints validation by sampling resulted in the distributions shown in Fig.
+5.
+Due to the non-linearity of the bid functions and the non-normal, non-symmetric marginals, the constraints with $\lambda$ values for normal distributions are no longer met perfectly.
+In principle, for each chance-constraint a local search for optimal $\lambda$ values is possible.
+Another option is the use of a distributionally robust $\lambda$, which however leads to conservative estimates as $\lambda_{1}^{rob},\lambda_{99}^{rob}=\mp\sqrt{\frac{1-0.01}{0.01}}=\mp 9.950$
+Figure 4: Sampled bid function distributions computed both with the bid functions and gPC expansions for the eight bids in the case of arbitrarily distributed marginals and polynomial bid functions.
+Figure 5: Sampled procured power distributions computed with the gPC expansion in the case of arbitrarily distributed marginals.
+For comparison the PDFs both considering and not considering correlation were sampled.
+The empirical $1^{st}$ and $99^{th}$ percentiles are indicated for verification of the chance-constraints.
 IV Conclusion
+In this paper, we developed a tractable framework for modeling correlated stochastic energy resources using Polynomial Chaos Expansion combined with Gaussian Copulas.
+The approach enables to reformulate chance constraints on weighted sums of random functions as convex constraints.
+Several ideas were introduced to improve computational efficiency, including transformations for quadrature integration of Gaussian Copulas as well as precomputing expectation values for parallelization and effective dimensionality reduction.
+The framework was validated on a numerical joint procurement problem with eight random variables and bids.
+First, using normally distributed marginals and linear bid functions, we showed that the reformulation correctly meets constraints for a correlated joint PDF.
+Then, for a more representative example with polynomial bid functions and Beta distributed marginals, a 21-dimensional orthonormal polynomial basis and corresponding PCE coefficients could be computed within two minutes.
+The optimization closely satisfies the chance constraints when using Gaussian quantile factors.
+Key limitations include the reliance of dimensionality reduction techniques on the dimensionality of the random functions as well as the need for quantile factors $\lambda$ in the exact reformulation of chance constraints.
+While $\lambda$ is known for normal distributions and conservative robust bounds are available, its exact value remains generally unknown for arbitrary distributions.
+Future work will focus on extending the framework to time-coupled problems, accounting for correlations across time steps.
+In addition, the formulation of joint chance constraints will be investigated.

### v3

@@
 I Introduction
-distributed energy resources are increasingly participating in ancillary service (AS) markets.
-Many of these resources.
-such as heat pumps, electric vehicles and photovoltaic (PV) installations have uncertain capabilities.
-To facilitate the participation of these DERs, some transmission system operators already allow bids for which the full power availability is only guaranteed with a certain probability [1].
-This has important consequences for the reliable procurement of reserve capacity.
-This raises new questions with respect to the amount of reserve that should be procured, as the availability of the procured reserve becomes a stochastic quantity.
-Recently, generalised polynomial chaos (gPC) has been used in power systems applications to model propagation of uncertain input variables.
+[…]
+DER are increasingly participating in ancillary services markets.
+Many of these resources, such as heat pumps, electric vehicles and photovoltaic (PV) installations have uncertain capabilities.
+To facilitate the participation of these distributed energy resources, some transmission system operators already allow bids for which the full power availability is only guaranteed with a certain probability [4].
+This raises new questions with respect to the amount of ASs that should be procured, as the full availability of the contracted ASs becomes uncertain.
+Existing works considering probabilistic reserve procurement either consider distributionally robust methods to account for uncertainties (e.g.
+[15] or [13]) or formulate chance constraints for the availability of different reserve bids, but do not account for the dependencies between the stochastic bid availabilities (e.g.
+[5]).
+Recently, generalised polynomial chaos (gPC) has been used in power systems applications to model the propagation of uncertain input variables.
 Advantages include the ability to model general probability distributions and represent stochastic variables through orthogonal polynomial expansions.
@@
 For Gaussian random variables or distributionally robust (DR) uncertainty, exact reformulations are found [11].
-For other distributions, [3] proposes an iterative tuning approach for the bounds in the moment-based reformulation in an attempt to recover the desired constraint satisfaction rate.
-A major limitation of these approaches, is that they consider the different random input variables to be independent.
+For other distributions, iterative tuning approaches for the bounds in the moment-based reformulation exist [7] attempting to recover the desired constraint satisfaction rate.
+A major limitation of these approaches is, that they consider the different random input variables to be independent.
+This can lead to underestimating the probability of violating certain thresholds.
 To allow modelling dependencies between random variables, [12] proposes to construct an orthonormal polynomial basis using a Gram-Schmidt (GS) procedure.
@@
 For linear dependencies, this can be simplified to the Nataf transform [8], but the authors of [6] show that it performs worse than their proposed GS procedure.
-Other approaches, so-called domination methods, approximate functions of dependent random variables with functions of independent random variables, but only show good performance when an accurate approximation can be obtained [2].
-None of the works above handling reserve procurement allow to assess the probability of violating the power reserve requirements when considering uncertain bids exhibiting dependencies.
-Additionally, the proposed approaches to extend gPC to dependent variables, lead to computational issues due to error propagation and computation time.
+Other approaches, so-called domination methods, approximate functions of dependent random variables with functions of independent random variables, but only show good performance when an accurate approximation of the orthogonal basis in terms of inner products with respect to the original joint distribution can be obtained [6].
+None of the works above handling reserve procurement allow to consider uncertain bids exhibiting dependencies.
+Additionally, the proposed approaches to extend gPC to dependent variables lead to computational issues due to error propagation and computation time.
 Therefore, this paper makes the following contributions:
@@
   3. 3.
-    A convex reformulation for probabilistic power reserve procurement considering dependent random variables.
+    A convex reformulation for probabilistic AS procurement considering dependent random variables.
 II Methodology
 II-A gPC with Dependent Random Variables
-Consider the random vector $\bm{\xi}=[\xi_{1},\ldots,\xi_{d}]^{\top}$ of dimension $d$ with joint probability density function (PDF) $p(\bm{\xi})$ and support $\Xi$.
+Consider the $d$-dimensional random vector $\bm{\xi}=[\xi_{1},\ldots,\xi_{d}]^{\top}$ with joint probability density function (PDF) $p(\bm{\xi})$ and support $\Xi$.
 As shown in previous works (e.g.
-[8]), a square-integrable random function $q(\bm{\xi})$ can be represented by a PCE
+[12]), a square-integrable random function $q(\bm{\xi})$ can be represented by a PCE.
 $q(\bm{\xi})\cong\sum_{l=0}^{L}a_{l}\psi_{l}(\bm{\xi}).$ (1)
-where the equality becomes exact as the arbitrary number of terms $L$ goes to infinity. The basis polynomials $\{\psi_{l}\}$ are orthonormal with respect to the inner product
-$\left\langle\psi_{m},\psi_{n}\right\rangle=\int_{\Xi}\psi_{m}(\bm{\xi})\,\psi_{n}(\bm{\xi})\,p(\bm{\xi})\,\mathrm{d}\bm{\xi}.$ (2)
-The expansion coefficients $\{a_{l}\}$ are obtained by projection onto the polynomial basis. Since this inner product corresponds to an expectation, the coefficients read
-$a_{l}=\left\langle q,\psi_{l}\right\rangle=\mathbb{E}\!\left[q(\bm{\xi})\,\psi_{l}(\bm{\xi})\right].$ (3)
+with the equality becoming exact as the arbitrary number of terms $L$ goes to infinity. The basis polynomials $\{\psi_{l}\}$ are orthonormal with respect to the inner product
+$\left\langle\psi_{m},\psi_{n}\right\rangle=\int_{\Xi}\psi_{m}(\bm{\xi})\,\psi_{n}(\bm{\xi})\,p(\bm{\xi})\,\mathrm{d}\bm{\xi}$ (2)
+and the expansion coefficients $\{a_{l}\}$ are obtained by projection onto the polynomial basis. Since this inner product corresponds to an expectation, the coefficients read
+$a_{l}=\left\langle q,\psi_{l}\right\rangle=\mathbb{E}\!\left[q(\bm{\xi})\,\psi_{l}(\bm{\xi})\right]$ (3)
+The moments of the output distribution of $q(\bm{\xi})$ can be recovered from the expansion coefficients. For instance, the mean $\mu$ and variance $\sigma$ are given by
+$\mu$ $=a_{0}$ (4)
+$\sigma$ $\cong\left\|[a_{1},\ldots,a_{L}]^{\top}\right\|_{2}$
 The main challenge lies in the construction of orthonormal polynomial bases with respect to an arbitrary joint PDF $p(\bm{\xi})$.
-For independent random variables, orthogonal polynomial families associated with the Askey scheme can be employed, since they are orthogonal with respect to their corresponding marginal distributions (e.g.
+For independent random variables, polynomial families associated with the Askey scheme can be employed, since they are orthogonal with respect to their corresponding marginal distributions (e.g.
 Hermite polynomials for Gaussian random variables).
@@
 all monomials with degree $\leq\nu$).
-As noted, a GS procedure can be used for the construction, but is numerically ill-conditioned [2].
+As noted in the introduction, a GS procedure can be used for the construction, but is numerically ill-conditioned [6].
 To address this, [14] provides an alternative algorithm based on a Whitening transformation, relying only in part on GS procedures.
-Finally, the moments of the output distribution of $q(\bm{\xi})$ can be recovered from the expansion coefficients. For instance the mean and variance are given by
-$\mu$ $=a_{0}$ (4)
-$\sigma$ $\cong\left\|[a_{1},\ldots,a_{L}]^{\top}\right\|_{2}$
 II-B Joint Probability using Copulas
 For practical modelling of statistically dependent random variables, the joint PDF $p(\bm{\xi})$ must be approximated in a tractable manner.
-This requires specifying both the marginal distributions of the individual components $f(\xi_{i})$ as well as their dependence structure.
+This requires specifying both the marginal distributions of the individual components $f_{i}(\xi_{i})$ as well as their dependence structure.
 Copulas provide a general and flexible framework for separating the modelling of marginal distributions from the modelling of statistical dependence.
@@
 $F(\bm{\xi})=C\!\left(F_{1}(\xi_{1}),\ldots,F_{d}(\xi_{d})\right).$ (5)
-Sklar’s Theorem states, that for every multivariate CDF there exists a unique Copula capturing the general dependence structure given by (5).
-In practice, one of several parametric Copula families is used.
-For the application considered in this paper no specific tail-dependence is expected.
-As a result, a Gaussian Copula is selected to maintain computational tractability and leverage its closed-form mathematical expressions.
+Sklar’s Theorem states that for every multivariate CDF there exists a Copula capturing the general dependence structure given by (5).
+In practice, one of the several parametric Copula families is used.
+For the application considered in this paper, a Gaussian Copula is selected.
+The Gaussian Copula captures dependence through a correlation matrix $\bm{\Sigma}$ containing the pairwise correlations between the random variables.
 Identifying $u_{i}=F_{i}(\xi_{i})$, it is defined as:
 $C_{\bm{\Sigma}}(\bm{u})=\Phi_{\bm{\Sigma}}(\Phi^{-1}(u_{1}),...,\Phi^{-1}(u_{d}))$ (6)
-where $\Phi_{\bm{\Sigma}}(\bm{z})$ is the multivariate normal CDF with correlation matrix $\bm{\Sigma}$ and $\Phi^{-1}(u_{i})$ the inverse normal CDF. The joint PDF is then expressed as
+where $\Phi_{\bm{\Sigma}}(\bm{z})$ is the multivariate normal CDF with correlation matrix $\bm{\Sigma}$ and $\Phi^{-1}(u_{i})$ the inverse normal CDF.
+This closed-form expression allows efficient scaling to higher dimensions while maintaining computational tractability.
+In contrast, Archimedean Copulas provide limited flexibility in higher dimensions due to their single-parameter structure, whereas Vine Copulas can represent complex and asymmetric dependence structures but rapidly become intractable as dimensionality increases.
+Using a Copula, with its density expressed as $c(\bm{u})$, the joint PDF can be formulated as
 $p(\bm{\xi})=c(\bm{u})*\prod_{i=1}^{d}f_{i}(\xi_{i})$ (7)
-where $c(\bm{u})$ is the Copula density. For a Gaussian Copula, the density is given by
-$c_{\bm{\Sigma}}(\bm{u})=\frac{\phi_{\bm{\Sigma}}\big(\Phi^{-1}(u_{1}),\ldots,\Phi^{-1}(u_{d})\big)}{\prod_{i=1}^{d}\phi\big(\Phi^{-1}(u_{i})\big)},$ (8)
-where $\phi_{\bm{\Sigma}}(\bm{x})$ is the multivariate normal PDF with correlation matrix $\bm{\Sigma}$} and $\phi(x_{i})$ the normal PDF.
+For a Gaussian Copula, the density is given by (8), where $\phi_{\bm{\Sigma}}(\bm{x})$ is the multivariate normal PDF with correlation matrix $\bm{\Sigma}$} and $\phi(x_{i})$ the normal PDF.
+$c_{\bm{\Sigma}}(\bm{u})=\frac{\phi_{\bm{\Sigma}}\big(\Phi^{-1}(u_{1}),\ldots,\Phi^{-1}(u_{d})\big)}{\prod_{i=1}^{d}\phi\big(\Phi^{-1}(u_{i})\big)}$ (8)
 II-C Quadrature Integration
 Both the construction of the orthonormal polynomial basis and the computation of the expansion coefficients require the evaluation of a large number of integrals of the form
-$\int_{\Xi}g(\bm{\xi})\cdot p(\bm{\xi})d\bm{\xi}$ (9)
+$\int_{\Xi}g(\bm{\xi})\,p(\bm{\xi})d\bm{\xi}$ (9)
 with $g(\bm{\xi})$ an arbitrary function.
@@
 Given the potentially high dimensionality of the integrals, Monte Carlo sampling is a natural choice.
-This however requires a large amount of samples to achieve sufficient accuracy and sampling errors may accumulate throughout the PCE process.
+This, however, requires a large amount of samples to achieve sufficient accuracy and sampling errors may accumulate throughout the PCE process.
 Moreover, the use of importance sampling is not straightforward in this context, since the integrands encountered vary across polynomial orders and random functions.
@@
 Let
-$u_{i}=F_{i}(\xi_{i})\in[0,1],\qquad i=1,\dots,d.$ (10)
-The Jacobian matrix of this transformation is diagonal with entries $\partial u_{i}/\partial\xi_{i}=f_{i}(\xi_{i})$ the marginal PDFs. Hence
-$d\bm{\xi}=\frac{d\bm{u}}{\prod_{i=1}^{d}f_{i}(\xi_{i})}.$ (11)
+$u_{i}=F_{i}(\xi_{i})\in[0,1]\qquad i=1,\dots,d$ (10)
+The Jacobian matrix of this transformation is diagonal with entries $\partial u_{i}/\partial\xi_{i}=f_{i}(\xi_{i})$ the marginal PDFs. Hence,
+$d\bm{\xi}=\frac{d\bm{u}}{\prod_{i=1}^{d}f_{i}(\xi_{i})}$ (11)
 Applying the change of variables and substituting $p(\bm{\xi})$ from (7) yields
-$\int_{\Xi}g(\bm{\xi})\cdot p(\bm{\xi})d\bm{\xi}=\int_{[0,1]^{d}}g\big(\bm{F}^{-1}(\bm{u})\big)\,c(\bm{u})\,d\bm{u}$ (12)
-where $\bm{F}^{-1}(\bm{u})$ is short-hand notation for the componentwise inverse marginal CDF.
-Evidently, this transformation is valid no matter the Copula used, in principle allowing any Copula family to be used for this quadrature integration scheme.
+$\int_{\Xi}g(\bm{\xi})\,p(\bm{\xi})d\bm{\xi}=\int_{[0,1]^{d}}g\big(\bm{F}^{-1}(\bm{u})\big)\,c(\bm{u})\,d\bm{u}$ (12)
+where $\bm{F}^{-1}(\bm{u})$ is short-hand notation for the component-wise inverse marginal CDF.
+Evidently, this transformation does not rely on the specific Copula used, in principle allowing any Copula family to be used for this quadrature integration scheme.
 However, the Copula density $c(\bm{u})$ is in general neither polynomial nor constant, meaning Gauss-Legendre quadrature will be inexact.
 For the specific case of Gaussian Copulas, consider a second transformation to a Gaussian latent space
-$z_{i}=\Phi^{-1}(u_{i})\in(-\infty,\infty),\qquad i=1,\dots,d$ (13)
-on which tensor-grid Gauss-Hermite quadrature is applicable. The Jacobian matrix is again diagonal with entries $\partial u_{i}/\partial z_{i}=\phi(z_{i})$, hence
+$z_{i}=\Phi^{-1}(u_{i})\in(-\infty,\infty)\qquad i=1,\dots,d$ (13)
+on which tensor-grid Gauss-Hermite quadrature is applicable. The Jacobian matrix is again diagonal with entries $\partial u_{i}/\partial z_{i}=\phi(z_{i})$, meaning
 $d\bm{u}=\Big(\prod_{i=1}^{d}\phi(z_{i})\Big)\,d\bm{z}$ (14)
 Applying the change of variables and substituting the Gaussian Copula density from (8) yields
-$\int_{\Xi}g(\bm{\xi})\cdot p(\bm{\xi})d\bm{\xi}=\int_{\mathbb{R}^{d}}g\big(\bm{F}^{-1}(\bm{\Phi}(\bm{z}))\big)\;\phi_{\bm{\Sigma}}(\bm{z})\,d\bm{z}$ (15)
-with $\bm{\Phi}(\bm{z})$ the componentwise normal CDF.
+$\int_{\Xi}g(\bm{\xi})\,p(\bm{\xi})d\bm{\xi}=\int_{\mathbb{R}^{d}}g\big(\bm{F}^{-1}(\bm{\Phi}(\bm{z}))\big)\;\phi_{\bm{\Sigma}}(\bm{z})\,d\bm{z}$ (15)
+with $\bm{\Phi}(\bm{z})$ the component-wise normal CDF.
 The result in (15) features a correlated normal PDF.
@@
 $\bm{y}_{n}=\bm{L}\bm{z}_{n}$ (16)
-allows to “colour“ the correlation onto the quadrature nodes resulting in
+allows to “color“ the correlation onto the quadrature nodes resulting in
 $\int_{\Xi}g(\bm{\xi})\,p(\bm{\xi})\,d\bm{x}\;\approx\;\sum_{n=1}^{N}w_{n}\cdot g(\bm{F}^{-1}(\Phi(\bm{Lz}_{n})))$ (17)
@@
 Due to this exponential scaling, dimensionality reduction strategies are necessary.
-Note that while the second transformation eliminated the non-polynomial Copula density $c(\bm{u})$, improving numerical accuracy, there is still no exactness guarantee.
-Gauss-Hermite quadrature is exact for polynomials up to degree 2k12k-1 weighted by a normal PDF, but $g(\bm{F}^{-1}(\bm{\Phi}(\bm{Lz})))$ does not necessarily behave like a polynomial of degree $\leq 2k-1$ in Gaussian latent space.
+Note that, even though the second transformation eliminates the non-polynomial Copula density $c(\bm{u})$ and improves numerical accuracy, there is still no exactness guarantee.
+Tensor-grid Gauss–Hermite quadrature is exact only for integrands that are polynomials up to degree 2k12k-1 in each variable weighted by a normal PDF, but $g(\bm{F}^{-1}(\bm{\Phi}(\bm{Lz})))$ does not necessarily behave like a polynomial in Gaussian latent space.
+For a comprehensive treatment of quadrature rules and their exactness properties, see classical texts such as [1].
 II-C1 Numerical Improvements
@@
 $\mathbb{E}[\xi_{1}\xi_{2}]=\int_{\Xi_{1,2}}\xi_{1}\xi_{2}\,p_{1,2}(\xi_{1},\xi_{2})\,d\bm{\xi}_{1,2}.$ (19)
-where $p_{1,2}(\xi_{1},\xi_{2})$ is a reduced joint PDF.
-For elliptic Copulas they have identical structure to the full joint PDF defined in (7), just with less input variables.
-For other Copulas, the reduced joint PDFS might need to be retrieved separately.
-Now, instead of directly applying an orthonormalization scheme, the monomial expectation values of all pairs of monomials in the chosen set can be evaluated beforehand.
+where $\Xi_{1,2}$ is the corresponding support and $p_{1,2}(\xi_{1},\xi_{2})$ the joint PDF of the appearing random variables.
+For elliptic Copulas, like the Gaussian Copula, such lower dimensional joint PDFs have identical structure to the full joint PDF defined in (7) and are therefore easy to recover.
+For other Copulas, they might need to be computed separately.
+Now, instead of directly applying an orthonormalization scheme, the expectation values of all pairwise products of monomials in the chosen set can be evaluated beforehand.
 When applying the orthonormalization scheme, the inner products can be expanded as in (18) and the expectation values substituted.
-Crucially, precomputing expectation values can be done in a parallel fashion and naturally limits the maximum integral dimension to be evaluated to $2\nu$.
-This because the highest dimensional expectation values to be evaluated are of the form $\mathbb{E}[\xi_{1}\xi_{2}\ldots\xi_{2\nu}]$.
-This approach decouples the integral dimension $d_{I}$ from the random variable dimension $d$.
+Crucially, precomputing expectation values can be done in a parallel fashion and naturally limits the maximum integral dimension to be evaluated to $2\nu$, as the highest dimensional expectation values to be evaluated are of the form $\mathbb{E}[\xi_{1}\xi_{2}\ldots\xi_{2\nu}]$.
+This decouples the integral dimensionality $d_{I}$ from the random variable dimension $d$.
 In the case where multiple random functions $q_{i}(\bm{\xi_{i}})$ only depending on subsets $\bm{\xi_{i}}$ of the random vector are expanded to the same polynomial basis, a similar approach can be used for the expansion coefficients.
-Here the expectation values $\mathbb{E}[q_{i}(\bm{\xi_{i}})],\ldots,\mathbb{E}[q_{i}(\bm{\xi_{i}})\xi_{d}^{\nu}]$ can be computed before with reduced dimension.
+Here the expectation values $\mathbb{E}[q_{i}(\bm{\xi_{i}})],\ldots,\mathbb{E}[q_{i}(\bm{\xi_{i}})\xi_{d}^{\nu}]$ can be computed beforehand with reduced dimension.
 Note, that the reduction in this case is dependent on the dimensionality of the random functions $q_{i}(\bm{\xi_{i}})$.
-Finally, the set of monomials from which the polynomial basis is constructed can be tuned.
+The implementation of the quadrature integration procedure using Gaussian Copulas is summarized in Algorithm 1.
+Finally, the set of monomials from which the polynomial basis is constructed can be reduced to limit the computational cost.
 For instance, assume the random functions $q_{i}(\bm{\xi_{i}})$ are models of PV power output at locations $i$ with $\bm{\xi_{i}}$ the local weather conditions like irradiation and temperature.
-Physically meaningless monomials such as the product of squares of the irradiations at two different locations can be dropped from the set without significantly affecting the expansion accuracy.
+Physically irrelevant monomials, such as higher-order products of the irradiations at different locations, can be dropped from the set without significantly affecting the PCE accuracy.
+This approach is conceptually similar to sparse PCE, which discards basis polynomials with negligible influence on the expansion output [2].
 III Use Case
@@
 To validate the formalism introduced in this paper, a simple cross-border joint procurement optimization with uncertain bids and chance constraints on the available power is solved.
-Formally, let $X$ and $Y$ be two zones sharing a common market with a set of bids $E_{i}(\bm{\xi})$ and costs $c_{i}$.
+Formally, let $X$ and $Y$ be two zones sharing a common AS market with a set of bids $E_{i}(\bm{\xi})$ and costs $\gamma_{i}$.
 Bids $i={1,\ldots,N_{X}}$ are located in zone $X$ and $i={N_{X}+1,\ldots,N_{X}+N_{Y}}$ are located zone Y.
 The decision variables $x_{i},y_{i}\in[0,1]$ determine the procured share of each bid.
-The objective is to minimize total cost, while both zones meet their reserve needs $R^{(X)},R^{(Y)}$ and stay below exchange limits $T_{X\to Y},T_{Y\to X}$ with $99\%$ probability.
-Expressed in standard form
-$\min_{\bm{x},\bm{y}}$ $\bm{c^{T}}(\bm{x}+\bm{y})$ (20)
+The objective is to minimize total cost, while both zones meet their reserve needs $R^{(X)},R^{(Y)}$ and stay below exchange limits of the interconnecting tie lines $T_{X\to Y},T_{Y\to X}$ with $99\%$ probability.
+The value of this probability threshold is inspired from Art 157.
+(h)-(i) in [3].
+Expressing the problem in standard form
+$\min_{\bm{x},\bm{y}}$ $\bm{\gamma}^{\top}(\bm{x}+\bm{y})$ (20)
 s.t. $0\leq\bm{x}+\bm{y}\leq 1$
@@
 For this, note that arbitrary weighted sums of gPC expansions result in a new gPC expansion
-$\sum_{j=1}^{Z}z_{j}E_{j}(\bm{\xi})$ $=\sum_{j=1}^{Z}z_{j}\sum_{l=0}^{L}a_{l}^{j}\psi_{l}(\bm{\xi})$ (21)
-$=\sum_{l=0}^{L}\bar{a}_{l}\psi_{l}(\bm{\xi})$
+Algorithm 1 Quadrature Integration with Dimension Reduction for a Gaussian Copula
+0: Integrand function $g(\bm{\xi})$, correlation matrix $\bm{\Sigma}\in\mathbb{R}^{d\times d}$, number of evaluation points per dimension $k$
+0: Numerical integral value II
+1: Determine the effective dimension $d_{I}$ of $g(\bm{\xi})$,
+2: Extract the submatrix $\bm{\Sigma}_{d_{I}}\in\mathbb{R}^{d_{I}\times d_{I}}$ corresponding to the relevant dimensions
+3: Compute the Cholesky factorization $\bm{LL}^{\top}=\bm{\Sigma}_{d_{I}}$
+4: Determine the $N=k^{d_{I}}$ tensor-grid Gauss-Hermite quadrature nodes $\{\bm{z_{n}}\}$ and weights $\{w_{n}\}$
+[…]
+6:for $n=1$ to NN do
+7:  $I\leftarrow I+w_{n}\cdot g(\bm{F}^{-1}(\bm{\Phi}(\bm{Lz_{n}})))$
+8:end for
+9:
+10:return II
 Figure 1: Illustration of the use case with four bids per zone. Two bids each are distributed across four locations with each two local random variables describing the temperature and irradiation.
+$\sum_{j=1}^{Z}z_{j}E_{j}(\bm{\xi})$ $=\sum_{j=1}^{Z}z_{j}\sum_{l=0}^{L}a_{l}^{j}\psi_{l}(\bm{\xi})=\sum_{l=0}^{L}\bar{a}_{l}\psi_{l}(\bm{\xi})$ (21)
 where $\bar{a}_{l}=\sum_{j=1}^{Z}z_{j}a_{l}^{j}$. Using the properties in (4), the mean and variance of such a weighted sum can be expressed as functions of the weights
@@
 For normal distributions $\lambda_{1},\lambda_{99}=\mp 2.326$.
-Applying this reformulation to all chance-constraints in problem (20) results in a convex optimization problem.
+Applying this reformulation to all chance-constraints in problem (20) results in the convex optimization problem (24).
 $\min_{\bm{x},\bm{y}}$ $\bm{c^{T}}(\bm{x}+\bm{y})$ (24)
@@
 $\sum_{i=N_{X}+1}^{N_{X}+N_{Y}}x_{i}a_{0}^{i}+\lambda_{99}\left\|(\bm{Ax})_{N_{X}+N_{Y}}^{N_{X}+1}\right\|_{2}\leq T_{Y\to X}$
-We solved a concrete example with $N_{X},N_{Y}=4$ bids per zone and $d=8$ random variables.
-In two situations.
-The bid functions are assumed to represent PV and HP capabilities.
-The random variables stand for the temperature and irradiation at four different locations with each two bids.
+We consider a concrete example with $N_{X},N_{Y}=4$ bids per zone and $d=8$ random variables and analyze two distinct cases.
+The bids are distributed across four locations, with two bids at each location depending on two random variables that represent local conditions.
 Fig.
-1 illustrates the setup.
+1 illustrates the use case.
 III-1 Normal marginals
-As $\lambda$ values are known exactly for normal distributions, the full process is first validated with arbitrary normally distributed marginals and linear bid functions.
-The correlation matrix $\bm{\Sigma}$ is fit from synthetic weather data of four different locations retrieved from renewables.ninja [9], however as there is no validation with real data, an arbitrary $\bm{\Sigma}$ could be used.
+As $\lambda$ values are known exactly for normal distributions, the full process is first validated with arbitrary normally distributed marginals and linear bid functions, together with an arbitrary correlation matrix $\bm{\Sigma}$.
 The polynomial basis is constructed from 9 monomials (all monomials of degree $\leq 1$).
@@
 Moreover, the gPC expansion accurately approximates the bid functions.
-The maximum numerical error of the expansion was found to be on the order of $10^{-7}\%$
-Figure 2: Sampled bid function distributions computed both with the bid functions and gPC expansions for the eight bids in the case of normally distributed marginals and linear bid functions.
-Problem (24) is then solved with parameters $R^{(X)}$, $R^{(Y)}$, $T_{X\to Y}$, $T_{Y\to X}=100$ and $c_{1-8}=1$ to obtain optimal $\bm{x},\bm{y}$.
-The chance-constraints in (20) are validated by sampling both from the Copula and from the marginals in an uncorrelated.
+The maximum numerical error of the expansion is found to be on the order of $10^{-7}\%$, making the two distributions indistinguishable.
+Problem (24) is solved with parameters $R^{(X)}$, $R^{(Y)}$, $T_{X\to Y}$, $T_{Y\to X}=100$ and $\gamma_{1-8}=1$ to get optimal $\bm{x},\bm{y}$.
+The chance-constraints in (20) are validated by sampling both from the Copula and from the marginal distributions assuming independence.
 The resulting distributions are shown in Fig.
 3.
-As expected, the optimization problem correctly satisfies the chance-constraints only for the correlated case.
-This results suggests, that the presented formalism works as intended.
+As expected, the optimization problem tightly satisfies the chance-constraints only for the dependent case.
+This result shows that the presented formalism works as intended.
+Figure 2: Sampled bid function distributions computed both with the bid functions and gPC expansions for the eight bids in the case of normally distributed marginals and linear bid functions.
 Figure 3: Sampled procured power distributions computed with the gPC expansion in the case of normally distributed marginals.
-For comparison the PDFs both considering and not considering correlation were sampled.
+For comparison the PDFs both considering and not considering dependence are sampled.
 The empirical $1^{st}$ and $99^{th}$ percentiles are indicated for verification of the chance-constraints.
 III-2 Arbitrary marginals
-Obviously, bid functions are in general not linear in the random variables and the marginals not normally distributed.
+In general, bid functions are not linear in the random variables and the marginals are not normally distributed.
 For example, solar irradiance uncertainty is frequently modelled using Beta distributions due to its bounded support.
-For validation in a more realistic setting, the bid functions are replaced by simple polynomial models of PV and HP capability and the marginals by representative Beta distributions.
-The polynomial basis was constructed from 21 physically relevant monomials.
-The computation of the basis and all coefficients was achieved in roughly 2 minutes total with a non-parallelised implementation and $k=15$ quadrature points per dimension.
-Validating the gPC expansions with sampling resulted in the bid power distributions shown in Fig.
+For validation in a more realistic setting, the bid functions are replaced by simple polynomials and the marginals by representative Beta distributions.
+The polynomial basis is constructed from 21 relevant monomials.
+The computation of the basis and all coefficients is achieved in roughly 2 minutes total with a non-parallelised implementation and $k=15$ quadrature points per dimension.
+Validating the gPC expansions with sampling results in the bid power distributions shown in Fig.
 4.
-The maximum expansion error was found to be on the order of $10^{-6}\%$
-Problem (24) was solved with parameters $R^{(X)}$, $R^{(Y)}=1000$, $T_{X\to Y}$, $T_{Y\to X}=500$ and $c_{1-8}=1$.
-The constraints validation by sampling resulted in the distributions shown in Fig.
+The maximum expansion error is found to be on the order of $10^{-6}\%$, again with the two distributions indistinguishable.
+Problem (24) is solved with parameters $R^{(X)}$, $R^{(Y)}=1000$, $T_{X\to Y}$, $T_{Y\to X}=500$ and $\gamma_{1-8}=1$.
+The validation of constraints satisfaction through sampling results in the distributions shown in Fig.
 5.
-Due to the non-linearity of the bid functions and the non-normal, non-symmetric marginals, the constraints with $\lambda$ values for normal distributions are no longer met perfectly.
-In principle, for each chance-constraint a local search for optimal $\lambda$ values is possible.
-Another option is the use of a distributionally robust $\lambda$, which however leads to conservative estimates as $\lambda_{1}^{rob},\lambda_{99}^{rob}=\mp\sqrt{\frac{1-0.01}{0.01}}=\mp 9.950$
+Due to the non-linearity of the bid functions and the non-normal, non-symmetric marginals, the constraints with $\lambda$ values for normal distributions are no longer met tightly.
+Note, that the overall computation time is dominated by the PCE process, while the choice of parameters has no noticeable impact, besides determining the feasibility of the problem.
 Figure 4: Sampled bid function distributions computed both with the bid functions and gPC expansions for the eight bids in the case of arbitrarily distributed marginals and polynomial bid functions.
 Figure 5: Sampled procured power distributions computed with the gPC expansion in the case of arbitrarily distributed marginals.
-For comparison the PDFs both considering and not considering correlation were sampled.
+For comparison the PDFs both considering and not considering dependence were sampled.
 The empirical $1^{st}$ and $99^{th}$ percentiles are indicated for verification of the chance-constraints.
 IV Conclusion
-In this paper, we developed a tractable framework for modeling correlated stochastic energy resources using Polynomial Chaos Expansion combined with Gaussian Copulas.
+In this paper, we develop a tractable framework for modelling correlated random variables using PCE combined with Gaussian Copulas.
 The approach enables to reformulate chance constraints on weighted sums of random functions as convex constraints.
-Several ideas were introduced to improve computational efficiency, including transformations for quadrature integration of Gaussian Copulas as well as precomputing expectation values for parallelization and effective dimensionality reduction.
-The framework was validated on a numerical joint procurement problem with eight random variables and bids.
-First, using normally distributed marginals and linear bid functions, we showed that the reformulation correctly meets constraints for a correlated joint PDF.
+Several ideas are introduced to improve computational efficiency, including transformations for quadrature integration of Gaussian Copulas as well as precomputing expectation values for parallelization and effective dimensionality reduction.
+The framework is validated on a numerical joint procurement problem with eight random variables and bids.
+First, using normally distributed marginals and linear bid functions, we show that the reformulation tightly meets constraints for a correlated joint PDF.
 Then, for a more representative example with polynomial bid functions and Beta distributed marginals, a 21-dimensional orthonormal polynomial basis and corresponding PCE coefficients could be computed within two minutes.
 The optimization closely satisfies the chance constraints when using Gaussian quantile factors.
+In addition to the methodological contributions, the proposed framework has relevance for system operators.
+For instance, TSOs engaged in interzonal AS provision or distribution system operators participating in reserve procurement can potentially benefit from the ability to model correlated stochastic resources in a tractable optimization setting.
 Key limitations include the reliance of dimensionality reduction techniques on the dimensionality of the random functions as well as the need for quantile factors $\lambda$ in the exact reformulation of chance constraints.